OpenStudy (anonymous):

I have a geometry question ... its simple but quite irritating ..... am unable to solve it the perimeter of a circle is 18 times more then the area of right angled triangle .the radius is 2 times height of the triangle and base is 6 times more then the radius of the circle. calculate the area of the circle. answer option : 88 , 616, 308,176, 693 -- one of the answer is correct

OpenStudy (mertsj):

Can we agree on these equations: \[2\pi r=18(\frac{1}{2}bh)=9bh\] \[r=2h, h=\frac{r}{2}\] \[b=r+6\] So \[2\pi r=9(r+6)(\frac{r}{2})\]

OpenStudy (mertsj):

Where am I going wrong?

OpenStudy (mertsj):

Oh I see. I didn't read carefully the base is 6r

OpenStudy (phi):

fix typos There are a lot of facts to write down: p= 2 pi r p= 18 A A = 0.5 b h b= 6r r= 2h and we want to find pi r^2

OpenStudy (mertsj):

So\[2\pi r=9(6r)(\frac{r}{2})\]

OpenStudy (mertsj):

Which produces a radius of

OpenStudy (mertsj):

\[r=\frac{2\pi}{27}\]

OpenStudy (mertsj):

And an area of something that is way too small.

OpenStudy (phi):

which does not give us anything like the choices provided

OpenStudy (mertsj):

Exactly. So either I made a calculation error or we are misinterpreting the problem.

OpenStudy (phi):

No, I got the same thing. Maybe the problem is interpreting base is 6 times more then the radius

OpenStudy (mertsj):

Or the perimeter is 18 times more than the area of the triangle.

OpenStudy (phi):

reverse engineer the problem, to find what the base should be?

OpenStudy (mertsj):

6 times more than the radius implies we should add something to the radius to get the base but what???

OpenStudy (phi):

ashish should get some details from his teacher about this one.

OpenStudy (mertsj):

I agree. I have wasted quite enough time on it.